# DALK: Dynamic Co-Augmentation of LLMs and KG to answer Alzheimer’s Disease Questions with Scientific Literature

### v2

@@ Appendix A Details of LLMs for KG
Appendix B Details of KG for LLMs
In this section, we provide detailed input and output for adopting KG to augment LLMs, including path-based and neighbor-based sub-graph sampling results (Table 11), self-aware knowledge retrieval (Table 12), describing sub-graphs with LLMs (Table 13) and inference with sampled knowledge (Table 14). The question we showcase here is “The area of the brain resistant to Neurofibrillary tangles of Alzheimer’s disease is: A. Visual association areas B. Entorhinal coex C. Temporal lobe D.Lateral geniculate body”, which same as the one we use in Section 5.4.
@@ C. Temporal lobe
Appendix C ADQA Benchmark
@@
Appendix D Further Experiment for RAG

### v3

@@ Appendix B Details of LLMs for KG
Appendix C Details of KG for LLMs
In this section, we provide detailed input and output for adopting KG to augment LLMs, including path-based and neighbor-based sub-graph sampling results (Table 14), self-aware knowledge retrieval (Table 15), describing sub-graphs with LLMs (Table 16) and inference with sampled knowledge (Table 17). The question we showcase here is “The area of the brain resistant to Neurofibrillary tangles of Alzheimer’s disease is: A. Visual association areas B. Entorhinal coex C. Temporal lobe D.Lateral geniculate body”, which is the same as the one we use in Section 5.4.
@@ C. Temporal lobe
Appendix D ADQA Benchmark
@@
Appendix E Further Experiment for RAG
@@
Appendix F Time Cost for Subgraphs Extraction
MedQA MedMCQA NMMLU QA4MRE AVG
AVG Length 107.4 23.8 342.9 17.6 122.9
Time Cost (s) 2.25 0.89 2.25 1.09 1.62
Table 21: Average lengths of questions and time costs to construct the two subgraphs (path-based subgraph and neighbor-based subgraph) for each dataset.